What are two consecutive odd whole numbers such that the sum of their squares is 130?
Answer by ankor@dixie-net.com(22740)

What are two consecutive odd whole numbers such that the sum of their squares is 130?
:
x^2 + (x+2)^2 = 130
FOIL
x^2 + x^2 + 4x + 4 = 130
:
2x^2 + 4x + 4 - 130 = 0
:
2x^2 + 4x - 126 = 0
Simplify, divide by 2
x^2 + 2x - 63 = 0
Factor
(x+9)(x-7) = 0
Two solutions
x = 7 and 9
and
x = -9 and -7
:
Check both solutions in the original equation